Main Menu Contents Grades Notes Evaueck Prin Info Course Contents » Exam #4 >> 12-8.problem Print Info The Parker Company hashow do I get to the answer?
Answer #1
Purchase:
Purchase cost 16200
Less: Present value of residual value 1568 =2000*0.784
Net Present value 14632
Rent
Initial cost 8000
Add: Present value of annual operating costs 2597 =600*4.329
Net Present value 10597
As net present value of renting is less than that of purchasing, it is better to rent
Difference in favor of Renting 4035 =14632-10597
Better to rent by $4035